# Multi-Keypoint Affordance Representation for Functional Dexterous Grasping

### v1

@@ Abstract
Functional dexterous grasping requires precise hand-object interaction, going beyond simple gripping. Existing affordance-based methods primarily predict coarse interaction regions and cannot directly constrain the grasping posture, leading to a disconnection between visual perception and manipulation. To address this issue, we propose a multi-keypoint affordance representation for functional dexterous grasping, which directly encodes task-driven grasp configurations by localizing functional contact points. Our method introduces Contact-guided Multi-Keypoint Affordance (CMKA), leveraging human grasping experience images for weak supervision combined with Large Vision Models for fine affordance feature extraction, achieving generalization while avoiding manual keypoint annotations. Additionally, we present a Keypoint-based Grasp matrix Transformation (KGT) method, ensuring spatial consistency between hand keypoints and object contact points, thus providing a direct link between visual perception and dexterous grasping actions. Experiments on public real-world FAH datasets, IsaacGym simulation, and challenging robotic tasks demonstrate that our method significantly improves affordance localization accuracy, grasp consistency, and generalization to unseen tools and tasks, bridging the gap between visual affordance learning and dexterous robotic manipulation. The source code and demo videos will be publicly available at https://github.com/PopeyePxx/MKA.
@@ I Introduction
Functional dexterous grasping is a key capability for robots to perform complex object manipulations based on human instructions. Unlike traditional simple grasping, it requires a robotic dexterous hand to generate diverse grasping postures and make contact with different object regions depending on the task. This involves intricate physical interactions between the fingers and the object. For instance, in the “Hold Drill” task, the robot’s five fingers must firmly grasp the drill head, while in the “Press Drill” task, the index finger presses the drill switch while the other four fingers stabilize the handle. Thus, how to infer task-relevant object contact regions and grasping postures from visual perception is a fundamental challenge in functional dexterous grasping.
@@
Figure 1: Comparison between existing affordance-based grasping methods and our proposed Multi-Keypoint Affordance representation. (a) Existing methods identify only a rough interaction region, leading to uncertain interaction poses. (b) Our method localizes multiple keypoints corresponding to dexterous hand joints, enabling a unique and constrained grasping posture.
@@
In the field of vision, affordance-based methods [1, 2, 3, 4, 5] have been widely explored to predict potential human interaction regions. Deep-learning-based approaches estimate heatmaps [4, 5] or segmentation masks [2, 3] to indicate feasible interaction areas. However, existing methods [6, 7] can only provide coarse region predictions given an image and a task. A rough affordance map cannot specify the exact interaction posture, leading to uncertainty in the grasping motion and insufficient constraints for functional dexterous grasping, as shown in Fig. 1(a). Therefore, how to find a novel visual representation that not only identifies task-relevant contact areas but also directly constrains the dexterous grasping posture, ensuring a well-defined interaction between the hand and the object, remains a challenging problem.
@@
Keypoint-based representations offer a potential solution by structuring high-dimensional visual data into a compact and interpretable form. Many studies [8, 9, 10, 11] have demonstrated the effectiveness of keypoint-based approaches in robotic manipulation, often decomposing grasping tasks into object and environment keypoints. For instance, KETO [10] defines three types of keypoints: grasp points, functional points, and operation points. SKP [11] directly defines five keypoints on the object’s surface to support parallel grasping. However, these methods exhibit limitations in their visual representation: either the keypoints are manually defined for specific tasks, limiting generalization to novel objects and tasks, or they rely heavily on simulated environments for training, reducing their real-world applicability. Additionally, many of these approaches require extensive manual annotations, further increasing data collection costs.
@@
To improve the generalization and applicability of keypoint-based representations, VRB [12] introduces a more flexible visual representation by learning contact points and motion trajectories from human operation videos, demonstrating enhanced performance in robotic manipulation tasks. However, this method relies on post-processing steps, and its visual representation remains indirect. More recent advances in Large Vision Models (LVM) have significantly improved object feature extraction. For instance, ReKep [13] leverages LVMs to automatically extract candidate keypoints, and then filters them using vision-language models, directly guiding robotic operations. This approach enhances task generalization and establishes a more direct link between vision and action.
@@
To address the challenges, we propose the Multi-KeyPoint Affordance representation for Functional Dexterous Grasping. By localizing multiple keypoints on the object and the hand, a unique dexterous grasping posture with clear constraints is determined. First, we introduce the Contact-guided Multi-Keypoint Affordance (CMKA) learning, which leverages LVMs (e.g., SAM [14] and DINOv2 [15]) for fine-grained affordance feature extraction. The CMKA supervises Egocentric (Ego)-view images using hand-object interaction regions in Exocentric (Exo)-view images as contact priors via CAM [16], guiding keypoint learning towards meaningful functional contact areas and eliminating the need for manual keypoint annotations. Then, we introduce the Keypoint-based Grasp matrix Transformation (KGT) method to ensure consistent mapping between hand and object keypoints. We observe that the wrist, functional fingers (index or thumb), and little finger MCP joints effectively reflect the relative contact posture between the hand and the object. The positional relationship of these three points forms a unique triangular structure, providing a direct connection between hand and object keypoints. We conduct comprehensive experiments to evaluate the proposed framework for multi-point affordance localization across 6 tasks and 18 tool shapes on the public FAH dataset [17], achieving a 45.35% improvement over the state-of-the-art method in the KLD metric. In both IsaacGym [18] and real robot experiments, we successfully establish the geometric constraint relationship between tool and hand keypoints.
@@
  • A novel multi-keypoint affordance representation is proposed, which constrains dexterous grasping postures through the geometric relationships of keypoints in the hand-object interaction region, directly establishing a link between vision and dexterous grasping actions.
@@ II Related Work
Figure 2: The key process of learning and connecting visual perception to functional dexterous grasping actions. The CMKA learning module learns from Exo images with human operation experience and transfers the knowledge to Ego images, locating three keypoints constrained by dexterous grasping. The KGT method maps the hand-object relative pose, calculating the dexterous hand’s rotation and translation parameters ($R$, $T$) to control the grasping task.
@@ II-A Object Representation for Dexterous Grasping
Grasping and manipulation are fundamental topics in robotics. Traditional methods [19, 20, 21] often rely on six Degrees of Freedom (6DoF) poses to represent objects for parallel gripper tasks. However, these methods are insufficient for dexterous grasping, which requires handling multiple contact points and complex interactions beyond simple position and orientation. Early methods such as rigid body modeling [22, 23] and template matching [24, 25] are task-specific and lack generalization, limiting their applicability to diverse tasks. Recent studies have focused on object structure-based grasp affordance representations, such as ContactDB [26], which annotates object-finger contact relationships; the method in [27], which maps contact points to finger regions and intent codes; and F2F [28], which uses knowledge graphs to associate functional object parts with functional fingers. While these methods improve task performance, they depend on ideal perception systems that assume precise segmentation or localization of functional regions—an assumption rarely achievable in real-world settings. To address these limitations, we propose a robust object representation method specifically designed for dexterous grasping, eliminating the need for idealized perception inputs and enabling reliable handling of complex interactions.
@@ II-B Keypoint Representation and Robotic Manipulation
Keypoint-based methods have been widely used in computer vision tasks such as face recognition [29, 30], human pose estimation [31], and tracking [32], where keypoints typically serve as low-level features or part-level object descriptors. In robotics, keypoints provide compact representations of the environment and objects. For example, KETO [10] defines three types of keypoints—grasp points, functional points, and operation points—to describe tasks, whereas SKP [11] defines five keypoints directly on the object surface to support parallel grasping. However, these methods are task-specific and struggle to generalize to new tasks. Recent advancements in large models have introduced generalizable representations for robotic manipulation, such as ReKep [13], which employs LVMs [15, 14] to extract candidate keypoints and vision-language models to filter task-relevant keypoints for direct operational guidance. However, ReKep [13] focuses on simple parallel gripper tasks and requires additional reasoning steps, making it unsuitable for dexterous manipulation. Inspired by human hand interactions [33], we propose a multi-keypoint representation based on the wrist, functional fingers, and the little finger. This design directly constrains dexterous grasping postures, providing effective and robust solutions for complex manipulation tasks.
@@ II-C Visual Affordance and Interaction
Visual affordance learning explores potential object regions for specific actions and is a key topic in robotic grasping. Early fully supervised methods [34, 35] relied on large-scale annotated datasets, which were both expensive and time-consuming to create. To reduce annotation costs, recent research has shifted toward weakly supervised methods, leveraging keypoints [36, 37] or image-level labels [4, 38, 39]. In this work, we utilize human interaction images to supervise Ego-view images through contact features, significantly reducing training data costs by leveraging existing resources. Existing affordance methods for robotic manipulation, such as VRB [12], learn contact points and trajectories from human operation videos, whereas Robo-ABC [40] generates hand-object contact datasets to enable zero-shot generalization. Similarly, GAT [7] proposes pixel-level affordance learning to capture precise regions. However, these methods often depend on post-processing steps and additional modules, and their affordance regions are typically coarse, failing to provide the fine-grained constraints required for dexterous grasping. To address these limitations, we propose a multi-region keypoint affordance learning approach that directly provides fine-grained constraints tailored for dexterous grasping tasks.
Figure 3: Framework of the proposed CMKA. The framework includes: (1) a LVM-based Multi-Scale Clustering (LMSC) module for extracting candidate keypoints; (2) Keypoint feature extraction from egocentric (Ego) view; (3) Contact geometry knowledge transfer from exocentric (Exo) view to Ego view.
@@ III Methodology
In this study, our goal is to develop a complete system that establishes a direct visual representation for functional dexterous grasping and achieves cross-task and cross-object generalization. As shown in Fig. 2, during the training process, the proposed Contact-guided Multi-Keypoint Affordance (CMKA) learning module acquires human operation experience from exocentric (Exo) images and transfers it to egocentric (Ego) images, identifying three keypoints constrained by dexterous grasping in the Ego image. The details of this process will be explained in Sec. III-A. We then use the Keypoint-based Grasp matrix Transformation (KGT) method to map and constrain the hand-object relative pose using these keypoints, calculating the rotation and translation parameters ($R$, $T$) of the dexterous hand, thereby enabling control of the grasping task. The relevant details will be further described in Sec. III-B. During the inference process, only Ego images are required as input, and the parameters learned by CMKA can be used to identify the three affordance keypoints of the object.
@@ III-A Contact-guided Multi-KeyPoint Affordances Learning
To identify the keypoint regions on the object surface where the fingers should make contact, robust fine-grained feature extraction is required. To achieve this, as shown in Fig. 3, we first employ LVM-based Multi-Scale Clustering (LMSC) to obtain candidate keypoints from different parts of the object surface (see Sec. III-A1). Next, we perform keypoint feature extraction from the Ego-view (see Sec. III-A2) and design a keypoint weighting learning mechanism, which computes a weighted score for each candidate keypoint and selects the top three keypoints. Then, a keypoint-guided feature extraction module is used to perform deep feature extraction on the selected regions in the Ego-view images. Finally, we leverage cues from the Exo-view for knowledge transfer of the contact geometry relationship (see Sec. III-A3), using learnable Class Activation Mapping (CAM) technology [16] to extract hand-object interaction features from the Exo-view images, and employ cosine similarity loss to supervise keypoint selection in the Ego-view images, ensuring that the selected keypoints are concentrated around the hand-object contact regions.
III-A1 LVM-based Multi-Scale Clustering Module
Figure 4: Candidate keypoint selection using the proposed LMSC module.
Inspired by the ReKep  [13], we propose an LMSC module, which aims to focus clustering on finger contact regions, as shown in Fig. 4.
First, we extract multi-level features from multiple intermediate layers of the DINOv2 model [15] to capture information from low to high levels. Specifically, we extract features from the first three layers, denoted as $F_{li}$ (where $i=1,2,3$) and apply linear transformations and normalization to each layer. Next, a learnable weight vector $\alpha_{i}$ is used to perform weighted fusion, where the weights are normalized using the softmax function. The final fused feature representation is obtained as:
$F=\sum_{i=1}^{3}\alpha_{i}F_{li}.$
Then, $F$ is passed through a multi-layer perceptron (MLP) to obtain the feature representation $f$. To further incorporate multi-scale information, we perform upsampling and downsampling on $f$ to obtain $f_{\uparrow}$ and $f_{\downarrow}$, respectively. The final multi-scale feature representation is obtained by summation:
$f^{\prime}=f+f_{\downarrow}+f_{\uparrow}.$
Meanwhile, recent vision foundation models like Segment Anything Model  [14] (SAM) have demonstrated strong capacity to produce robust zero-shot segmentation in real-world scenes. we apply SAM  [14] to the Ego-view image to obtain multi-region masks $M_{i}$ and perform region-wise clustering on the multi-scale feature representation $f^{\prime}$ using the non-background masks. Specifically, we first apply PCA to reduce the dimensionality of each region’s features, obtaining the reduced feature representation $X_{\text{PCA}}\in\mathbb{R}^{n\times k}$, where $n$ is the number of pixels in the region and $k$ is the reduced dimension. Then, we perform K-means clustering on the reduced features to obtain multiple candidate keypoints $K_{cn}$, selecting the center of each cluster as the final candidate keypoint. The clustering aims to minimize the distance between samples and cluster centers:
$K_{cn}=\arg\min_{\mathbf{C}}\sum_{i=1}^{n}\|X_{\text{PCA}}^{(i)}-C_{j}\|^{2},$
where $K_{cn}$ represents the $n$-th candidate keypoint, $C_{j}$ is the center of the $j$-th cluster, and $X_{\text{PCA}}^{(i)}$ is the feature of the $i$-th sample in the region. Finally, we select the center of each cluster as the candidate keypoint. If no valid candidates are found, the pixel closest to the object centroid is chosen as a fallback keypoint.
III-A2 Keypoint Feature Extraction from Egocentric View

### v2

@@ Abstract
Functional dexterous grasping requires precise hand-object interaction, going beyond simple gripping. Existing affordance-based methods primarily predict coarse interaction regions and cannot directly constrain the grasping posture, leading to a disconnection between visual perception and manipulation. To address this issue, we propose a multi-keypoint affordance representation for functional dexterous grasping, which directly encodes task-driven grasp configurations by localizing functional contact points. Our method introduces Contact-guided Multi-Keypoint Affordance (CMKA), leveraging human grasping experience images for weak supervision combined with Large Vision Models for fine affordance feature extraction, achieving generalization while avoiding manual keypoint annotations. Additionally, we present a Keypoint-based Grasp matrix Transformation (KGT) method, ensuring spatial consistency between hand keypoints and object contact points, thus providing a direct link between visual perception and dexterous grasping actions. Experiments on public real-world FAH datasets, IsaacGym simulation, and challenging robotic tasks demonstrate that our method significantly improves affordance localization accuracy, grasp consistency, and generalization to unseen tools and tasks, bridging the gap between visual affordance learning and dexterous robotic manipulation. The source code and demo videos are publicly available at https://github.com/PopeyePxx/MKA.
@@ I Introduction
Functional dexterous grasping enables robots to execute complex object manipulations from human instructions. Unlike simple grasping, it requires a dexterous hand to adapt grasp postures and contact different object regions according to the task, involving fine physical interactions between the fingers and the object. For example, in “Hold Drill”, all five fingers firmly grasp the drill head, whereas in “Press Drill”, the index finger presses the switch while the others stabilize the handle. The core challenge lies in inferring task-relevant contact regions and grasp postures from visual perception.
@@
Figure 1: Comparison between existing affordance-based grasping methods and our proposed Multi-Keypoint Affordance representation. (a) Existing methods identify only a rough interaction region, leading to uncertain interaction poses. (b) Our method localizes multiple keypoints corresponding to dexterous hand joints, enabling a precise and constrained grasping posture.
@@
In the field of vision, affordance-based methods [1, 2, 3, 4, 5] have been widely explored to predict potential human interaction regions. Among them, deep-learning-based approaches estimate heatmaps [4, 5] or segmentation masks [2, 3] to indicate feasible interaction areas. However, existing methods [6, 7] can only provide coarse region predictions given an image and a task. A rough affordance map cannot specify the exact interaction posture, leading to uncertainty in the grasping motion and insufficient constraints for functional dexterous grasping, as shown in Fig. 1(a). Therefore, how to find a novel visual representation that not only identifies task-relevant contact areas but also directly constrains the dexterous grasping posture, ensuring a well-defined interaction between the hand and the object, remains a challenging problem.
@@
Keypoint-based representations offer a potential solution by structuring high-dimensional visual data into a compact and interpretable form. Prior work [8, 9, 10, 11] often decomposes grasping into object and environment keypoints. For example, KETO [10] defines grasp, functional, and operation points, while SKP [11] specifies five surface keypoints for parallel grasping. However, these approaches suffer from limited generalization—keypoints are either manually designed for specific tasks, rely heavily on simulation, or require extensive manual annotations, raising data collection costs.
@@
To address these issues, VRB [12] learns contact points and motion trajectories from human operation videos, improving generalization and applicability, though it still depends on post-processing and yields indirect visual representations. Recent advances in Large Vision Models (LVMs) enhance object feature extraction. For instance, ReKep [13] automatically detects candidate keypoints via LVMs, filters them with vision-language models, and directly guides robotic operations, improving task generalization and linking vision more directly to action.
@@
To address the challenges, we propose the Multi-KeyPoint Affordance representation for Functional Dexterous Grasping. By localizing multiple keypoints on the object and the hand, a unique dexterous grasping posture with clear constraints is determined. First, we introduce the Contact-guided Multi-Keypoint Affordance (CMKA) learning, which leverages LVMs for fine-grained affordance feature extraction. The CMKA supervises Egocentric images using hand-object interaction regions in Exocentric images as contact priors via CAM [14], guiding keypoint learning towards meaningful functional contact areas and eliminating the need for manual keypoint annotations. Unlike existing methods that only predict contact regions or keypoint locations without providing actionable grasp execution, we propose the Keypoint-based Grasp matrix Transformation (KGT) to bridge perception and control by explicitly computing the relative pose between the hand and the object. Specifically, KGT leverages the geometric relationship among three semantically meaningful keypoints-the wrist, functional finger (index or thumb), and little finger MCP joints-which form a unique triangular structure capturing the relative contact posture. Experiments across $6$ tasks and $18$ tool shapes on the public FAH dataset [15], achieving an improvement of $45.35\%$ over the state-of-the-art method in the KLD metric. In both IsaacGym [16] and real robot experiments, we successfully establish the geometric constraint relationship between tool and hand keypoints.
@@
  • A multi-keypoint affordance representation is proposed, which constrains dexterous grasping postures through the geometric relationships of keypoints in the hand-object interaction region, directly establishing a link between vision and dexterous grasping actions.
@@ II Related Work
Figure 2: The key process of learning and connecting visual perception to functional dexterous grasping actions. The Contact-guided Multi-Keypoint Affordance (CMKA) module learns from human demonstrations in Exo images and transfers this knowledge to Ego views, predicting three keypoints constrained by functional grasp priors. These keypoints are then used by the Keypoint-based Grasp matrix Transformation (KGT) module to compute the hand-object relative pose (($R$, $T$) to control the grasping task.
@@ II-A Object Representation for Dexterous Grasping
Grasping and manipulation are fundamental topics in robotics. Traditional methods [17, 18, 19] often rely on six Degrees of Freedom (6DoF) poses to represent objects for parallel gripper tasks. However, these methods are insufficient for dexterous grasping, which requires handling precise contact points on functional regions and complex interactions, going beyond simple object poses that only describe overall position and orientation. Early methods such as rigid body modeling [20, 21] and template matching [22, 23] are task-specific and lack generalization, limiting their applicability to diverse tasks. Recent studies have focused on object structure-based grasp affordance representations, such as ContactDB [24], which annotates object-finger contact relationships; the method in [25], which maps contact points to finger regions and intent codes; and F2F [26], which uses knowledge graphs to associate functional object parts with functional fingers. While these methods improve task performance, they depend on ideal perception systems that assume precise segmentation or localization of functional regions—an assumption rarely achievable in real-world settings. In contrast, we propose an object representation specifically designed for dexterous grasping. The object is abstracted as a set of semantically meaningful keypoints on its surface. This structured representation captures functional regions and supports executable grasp synthesis. It provides a bridging link between affordance perception and dexterous grasping control, eliminating the need for idealized perception inputs required by the existing methods [25, 26].
@@ II-B Keypoint Representation and Robotic Manipulation
Keypoint-based methods have been widely applied in computer vision tasks such as face recognition [27, 28], human pose estimation [29], and tracking [30], where keypoints typically serve as low-level features or part-level object descriptors. In robotics, keypoints provide compact representations of the environment and objects. For example, KETO [10] and SKP [11] define different types or fixed numbers of keypoints to describe specific tasks, but these methods often lack generalizability across tasks. Recently, ReKep [13] introduced a more generalizable manipulation framework by leveraging Large Visual Models (LVMs) [31, 32] to extract candidate keypoints and vision-language models to filter task-relevant keypoints for operational guidance. However, ReKep [13] focuses on simple parallel gripper tasks and requires additional reasoning steps, making it unsuitable for dexterous manipulation. Inspired by human hand interactions [33], we propose a multi-keypoint representation based on the wrist, functional fingers, and the little finger. This design directly constrains dexterous grasping postures, providing effective and robust solutions for complex manipulation tasks.
@@ II-C Visual Affordance and Interaction
Visual affordance learning explores potential object regions for specific actions and is a key topic in robotic grasping. Early fully supervised methods [34, 35] relied on large-scale annotated datasets, which were both expensive and time-consuming to create. To reduce annotation costs, recent research has shifted toward weakly supervised methods, leveraging keypoints [36, 37] or image-level labels [4, 38, 39]. In this work, we utilize human interaction images to supervise Ego-view images through contact features, significantly reducing training data costs by leveraging existing resources. Existing affordance methods for robotic manipulation, such as VRB [12], learn contact points and trajectories from human operation videos, whereas Robo-ABC [40] generates hand-object contact datasets to enable zero-shot generalization. Similarly, GAT [7] proposes pixel-level affordance learning to capture precise regions. However, those methods often rely on post-processing and additional modules, with coarse affordance regions that lack the fine-grained constraints needed for dexterous grasping.
Figure 3: Framework of the proposed CMKA: (1) LVM-based Multi-Scale Clustering for candidate keypoint extraction; (2) Affordance-aware keypoint weighting and feature extraction from Ego view; (3) Contact geometry knowledge transfer from Exo to Ego view.
@@ III Methodology
In this work, we propose a complete framework that establishes a direct visual representation for functional dexterous grasping with cross-task and cross-object generalization. As illustrated in Fig. 2, the proposed Contact-guided Multi-Keypoint Affordance (CMKA) module learns from human operation experience in exocentric (Exo) images and transfers this knowledge to egocentric (Ego) images, localizing three keypoints constrained by functional dexterous grasping (see Sec. III-A). Using these keypoints, the Keypoint-based Grasp matrix Transformation (KGT) method computes the hand-object relative pose, obtaining the rotation and translation parameters ($R$, $T$) required for grasp execution (see Sec. III-B). During inference, the system requires only an Ego image and the affordance class to predict the three functional keypoints on the object.
@@ III-A Contact-guided Multi-KeyPoint Affordances Learning
To identify the keypoint regions on the object surface where the fingers should make contact, robust fine-grained feature extraction is required. To achieve this, we first extract multi-level visual features from the Ego and Exo view images using the large vision model DINOv2 [31], which serves as the base visual representation for the following stages. As illustrated in the blue-shaded region of Fig. 3, we then apply LVM-based Multi-Scale Clustering (LMSC) to extract candidate keypoints from different parts of the object surface based on these features (see Sec. III-A1). In the green-shaded region, we design an affordance-aware keypoint weighting and feature extraction mechanism, where weights based on affordance class (e.g., “Press”) are assigned to candidate keypoints, guiding both the selection of the most relevant keypoints and their feature extraction from the Ego view (see Sec. III-A2). In the yellow-shaded region, we leverage the hand-object interaction feature extraction capability of the weakly-supervised LOCATE framework [38] for contact geometry knowledge transfer from Exo to Ego view. A cosine similarity loss is applied to supervise the keypoints selection, ensuring that the selected keypoints are concentrated around meaningful hand-object contact regions (see Sec. III-A3).
III-A1 LVM-based Multi-Scale Clustering Module (LMSC)
Figure 4: Candidate keypoint generation using the LMSC module.
Inspired by ReKep [13], we propose the LMSC module to improve keypoints selection in ego-view images, as illustrated in Fig. 4. Given an input image, we first apply the Segment Anything Model (SAM) [32] to generate a set of region masks $\{M_{i}\}_{i=1}^{S}$.
To improve the perception of fine-grained structural details and contact point features in the object, we then apply a Hierarchical Feature Fusion (HFF) mechanism to integrate multi-level and multi-scale features. Specifically, we extract features $F_{lm}$ (m=1,2,3m{=}1,2,3) from the last three layers of DINOv2. We then perform the fusion:
$F=\sum_{m=1}^{3}\alpha_{m}\cdot\varnothing(F_{lm}),$ (1)
where $\varnothing(\cdot)$ denotes a linear layer followed by normalization, and $\alpha_{m}$ are learnable weights. The fused feature $F$ is further processed by an MLP to obtain the dense feature $f$.
To capture multi-scale context, we generate bilinearly aligned upsampled features $f_{\uparrow}$ and downsampled features $f_{\downarrow}$ versions of $f$, and sum them with the original to obtain $f_{\text{ms}}$:
$f_{\text{ms}}=f+f_{\uparrow}+f_{\downarrow}.$ (2)
Then, we apply Principal Component Analysis (PCA) to $f_{\text{ms}}$ to obtain a reduced feature map $f_{\text{ms}}^{\text{PCA}}$, from which we extract region-wise features $X_{i}^{\text{PCA}}$.
Lastly, we introduce a two-step Cluster module, as illustrated in Fig. 4, to extract candidate keypoints from $X_{i}^{\text{PCA}}$. Specifically, we first apply K-means clustering on each $X_{i}^{\text{PCA}}$ to obtain $J$ cluster centers per region, resulting in a total of $N{=}S{\times}J$ cluster centers across all $S$ segmented regions:
$\{C_{n}\}_{n=1}^{N}=\{C_{ij}\}_{i=1,j=1}^{S,J}=\text{K-means}(\{X_{i}^{\text{PCA}}\}_{i=1}^{S}).$ (3)
These cluster centers are then mapped back to the pixel space by selecting the nearest feature vector from $X_{i}^{\text{PCA}}$, yielding the candidate keypoints $\{K_{n}\}_{n=1}^{N}$:
$K_{n}=\arg\min_{x\in X_{i}^{\text{PCA}}}\|x-C_{n}\|^{2}.$ (4)
If clustering is not applicable (e.g., due to insufficient pixels), the geometric centroid of $M_{i}$ is used as a fallback.
III-A2 Affordance-aware Keypoint Feature Extraction from Ego View